OpenStudy (anonymous):

could you guys tell me if this right? in terms of the format the question is Let f(x) = x**8 and g(x) = 6 x - 7. d/dx F(X)= 8x^7 d/dx g(x)= 6 but my problem is d/dx [f(g(x))] = (6(8x^7))?

OpenStudy (ash2326):

First find f(g(x) and then differentiate

OpenStudy (anonymous):

im not sure about (6(8x^7))

OpenStudy (ash2326):

Ok, I'll write f(g(x) for you then you find \(\frac{d}{dx} (f(g(x))\) \[f(g(x))={(g(x))}^8\] We get \[f(g(x))=(6x-7)^8\] Can you do now?

OpenStudy (ash2326):

f(g(x) is the composite function

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

becuase how are u getting to the 8th? you have even used the power rule yet

OpenStudy (ash2326):

I haven't differentiated it yet. Now you have to differentiate with respect to x and that will be your answer

OpenStudy (anonymous):

i know so u get 8 (6x-7)^7

OpenStudy (ash2326):

Yeah, That's correct. :D

OpenStudy (ash2326):

No you have to multiply it by 6 also \[\frac{d}{dx}(6x-7)=6\]

OpenStudy (anonymous):

ok but they only wanted me to find f(g(x))

OpenStudy (ash2326):

yeah , when you differentiate (6x-7)^8 we get \[7\times (6x-7)^7 \times 6\]

OpenStudy (anonymous):

ok thanx